# GiVE: Guiding Visual Encoder to Perceive Overlooked Information

### v1

@@ Abstract
Introduction
The multimodal large language models (MLLMs) (Zhu et al. 2023; Liu et al. 2023b; Ye et al. 2023; Li et al. 2023) have significantly advanced general artificial intelligence by demonstrating exceptional capabilities in generation and inference across diverse real-world applications, such as text-to-video generation (Blattmann et al. 2023), visual question answering (Singh et al. 2019), and embodied robotics (Mu et al. 2023). To optimize performance while managing costs, a prevalent MLLM architecture integrates a visual encoder with the large language model (LLM). This architecture encoder embeds non-text data, transforming these embeddings into vectors comprehensible to the LLM through a mapping mechanism. Although prior research (Wang et al. 2023; Liu et al. 2023a) has underscored the efficacy of this architecture, the cross-modal perception abilities of MLLMs heavily relies on the visual encoder, with the quality of encoded image embeddings directly influencing the response quality of the MLLMs.
An image encoder is a specialized visual encoder designed to map high-dimensional image data to a lower-dimensional feature space. These encoders can be broadly categorized based on their pre-training tasks into two main types: reconstruction-based and cross-modal contrastive learning-based encoders.
  • Image encoding models trained with reconstruction tasks (van den Oord, Vinyals, and Kavukcuoglu 2017; Huang, Fu, and Bors 2023) are proficient in capturing comprehensive image details. However, these models lack semantic alignment with text during training, which complicates the LLM’s ability to interpret image embeddings (Yin et al. 2023). Consequently, such encoders are infrequently utilized in MLLMs.
  • Vision transformer (ViT) models trained with image-text contrastive learning (Radford et al. 2021; Jia et al. 2021) generally align effectively with LLMs but face an implicit “ignore” problem, limiting their expressive capability. This limitation arises because different modalities convey distinct types of information. For instance, an image may feature multiple objects with unique attributes, such as texture, color, spatial location, and potential interactions. In contrast, abstract text typically highlights only the most salient objects and provides limited descriptions of other visual elements. ViTs trained for image-text matching tend to focus on the salient regions of the image that correspond to the text, thereby overlooking secondary elements like the background. As shown in Table 13 of the Appendix, MLLMs using such visual encoders exhibit diminished response quality when users inquire about non-salient objects.
In summary, for effective integration with LLMs, MLLMs require an image encoder that is both (1) semantically aligned with text during training and (2) capable of flexible attention to prevent the omission of relevant visual features.
@@ Introduction
To address these challenges, we propose this Guiding Visual Encoder to Perceive Overlooked Information (GiVE) approach, which aims to guide the visual encoder in adaptively adjusting its attention to well capture overlooked information. In this approach, we introduce a novel Attention-Guided Adapter (AG-Adapter) module that enhances the representation ability of the visual encoder by aligning the visual representations with abstract semantics. This module also functions as a plug-in for generalizing abstract semantics, enabling it to more effectively address user queries.
To tackle the above limitations in detail, GiVE incorporates another innovative module: Object-focused Visual Semantic Learning. This module employs three distinct model loss terms: (i) an Object-focused Image-Text Contrast (OITC) loss, which encourages the model to generate distinct embeddings for varied instructions, ensuring consideration of both salient and non-salient objects; (ii) an Object-focused Image-Image Contrast (OIIC) loss, which improves the accuracy of object retrieval by enabling the model to learn common features among in-class objects, thus enhancing concept generalization ability; and (iii) an Object-focused Image Discrimination (OID) loss, which improves the comprehensiveness of object retrieval by facilitating the model in identifying specific objects and recognizing potential correlations among objects, thereby preventing the omission of objects. As illustrated in Figure 1, GiVE enables visual encoders to generate image embeddings that are rich in targeted information.
@@
  • We propose an innovative GiVE approach to give attetion to overlooked information, enabling dynamic adjustment of focus by modifying textual instructions. This capability allows for flexible and adaptive attention to various objects within the image.
@@
Related Work
Pre-training Visual Encoders
Visual encoding models can be categorized based on their pre-training tasks into image reconstruction models and image-text contrast models. (i) Image reconstruction models, such as MAE (He et al. 2022), DeepMIM (Ren et al. 2023), and MIRL (Huang, Fu, and Bors 2023), use masked image modeling similar to language modeling. Other models, like VQ-VAE (van den Oord, Vinyals, and Kavukcuoglu 2017), VQGAN (Esser, Rombach, and Ommer 2021), and HQ-VAE (You et al. 2022), implement image reconstruction using autoencoder architectures. (ii) Models like CLIP (Radford et al. 2021), ALIGN (Jia et al. 2021), and EVA-CLIP (Sun et al. 2023) utilize cross-modal contrastive learning to align textual and visual information into a unified semantic space. Consequently, such pre-aligned encoders are more readily aligned with LLMs through alignment pre-training (Yin et al. 2023). In contrast, reconstruction-trained models lack this text correlation due to their independent training process. As detailed in Table 5 in the Appendix, 90% of the MLLMs we investigated use image-text contrast encoders.
Discussion. Our study preserves the semantic alignment capabilities of visual encoders while minimizing the omission of effective information by using semantic instructions.
@@
Visual Perception of MLLMs
Despite some pioneering works mitigating the negative effects of visual ignorance, they still have limitations. For instance, InstructBLIP (Dai et al. 2023) injects textual instructions to guide the post-processing process with a lightweight Q-Former to translate the image embeddings. However, this downstream secondary coding only reinforces indication-related information in the image embeddings and does not recover useful features lost in the upstream image encoder. Similarly, LLaVA-NeXT (Liu et al. 2024) encodes the image multiple times after segmentation, allowing each subimage to have its own salient object, thus reducing the visual feature loss rate. However, this exponentially increases token occupancy and encoding time. Additionally, BLIVA (Hu et al. 2024) combines the mapper of BLIP-2 and LLaVA to enhance the LLM’s understanding from more translation perspectives, but since both mappers share the same image encoder, this method does not fundamentally address the loss of useful information during visual encoding.
Dicussion. Overlooking the inherent imperfections of visual encoders and attempting to rectify them through alternative means is frequently inadequate. Our objective is to confront these shortcomings directly within the visual encoder, thereby minimizing the accumulation of errors from the outset of the encoding process.
Methodology
The proposed approach takes a text encoder to encode both captions and instructions, and a visual encoder for visual content. An AG-Adapter module is inserted into each stacked visual encoder as the tunable parameters to align cross-modal data. The AG-Adapter is trained by the designed Object-focused Visual Semantic Learning module. Details of model architecture, AG-Adapter, and object-focused visual semantic learning module are presented in the following subsections. The overall architecture is depicted in Figure 2.
Model Architecture
@@
The proposed model includes a visual encoder $\Phi_{I}(\cdot,\cdot)$ and a text encoder $\Phi_{T}(\cdot)$ to respectively encode visual and textual content. It accepts image-text-object triplets $\{\boldsymbol{x}^{I},\boldsymbol{x}^{T},\boldsymbol{x}^{O}\}$ as input, where $\boldsymbol{x}^{I}\in\mathcal{I}$ is an image, $\boldsymbol{x}^{T}\in\mathcal{T}$ is a text, and $\boldsymbol{x}^{O}\in\mathcal{O}$ is an indicative text denoting the target object, such as “person”. The model then extracts conditional image and text features $(\boldsymbol{y}^{I|O},\boldsymbol{y}^{T})$ using paired encoders. When extracting conditional image features, the instruction feature $\boldsymbol{y}^{O}$ is fused with the visual data stream within the AG-Adapter module. Formally,
$\boldsymbol{y}^{T}=\Phi_{T}(\boldsymbol{x}^{T}),\ \boldsymbol{y}^{O}=\Phi_{T}(\boldsymbol{x}^{O}),\ \boldsymbol{y}^{I|O}=\Phi_{I}(\boldsymbol{x}^{I},\boldsymbol{y}^{O}),$ (1)
@@
where $\boldsymbol{y}^{T}\in\mathbb{R}^{d}$ is text feature, $\boldsymbol{y}^{O}\in\mathbb{R}^{d}$ is instruction feature, and $\boldsymbol{y}^{I|O}$ is conditional image feature, i.e. the output of the visual encoder integrated with the AG-Adapter. The AG-Adapter is trained using our designed Object-focused Visual Semantic Learning component containing three loss terms: OITC, OIIC and OID. Note that during the training phase, the loss is calculated based on the output embeddings of both encoders. During the inference phase, the text encoder is retained to serialize the textual instructions.
@@
Figure 3: Learning objectives illustration. The image and text encoders jointly compute three losses. (a) For Object-focused Image-Text Contrast, the paired text and image should not only correspond to each other but also correspond to the same semantic object. (b) Object-focused Image-Image Contrast requires the model to predict pairs of image features that contain the same semantic object. (c) Object-focused Image Discrimination determines whether a specific object exists or not in the image. The text instructions, such as “bike”, are pre-integrated into the prompt template designated as “a photo of {object}”.
Attention-guided Adapter
@@
The proposed AG-Adapter module, as highlighted in green rectangle in Figure 2, is a simple yet effective plug-in that interweaves semantic directives with visual cues, enabling visual model to perceive queried objects. The AG-Adapter is inspired by Latent Diffusion Model (LDM) (Rombach et al. 2022) which enhances the alignment of visual representations with abstract semantics by conditioning on text representations. In particular, the AG-Adapter is incorporated into the pre-training feature extraction layers of the $\Phi_{I}(\cdot,\cdot)$, with only the inserted layer undergoing the training process.
@@
$\boldsymbol{f}^{V|O}=\varphi(\boldsymbol{y}^{O},\boldsymbol{f}^{V}),\vspace{-1mm}$ (2)
where $\boldsymbol{y}^{O}$ is instruction feature derived from user queries and $\boldsymbol{f}^{V}$ refers to the feature of the visual data flow. The $\boldsymbol{f}^{V}\in\mathbb{R}^{(M+K)\times d}$ is composed of two types of tokens: image tokens $\boldsymbol{f}^{I}\in\mathbb{R}^{M\times d}$ and other tokens $\boldsymbol{f}^{H}\in\mathbb{R}^{K\times d}$. The image tokens are derived directly from the input image, while the role of other tokens depends on the baseline model. In the case of GroupViT (Xu et al. 2022), other tokens are group tokens; in contrast, in CLIP (Radford et al. 2021), they are absent, and $K=0$. Within the AG-Adapter, the dual-layer MLP serves to bridge the gap between text and image representations, namely,
$\boldsymbol{f}^{O}=\text{MLP}(\boldsymbol{y}^{O}),\vspace{-1mm}$ (3)
@@
$\vspace{-1mm}\hat{\boldsymbol{f}}^{I}=\text{Norm}(\boldsymbol{f}^{I}),\vspace{-1mm}$ (4)
where $\text{Norm}(\cdot)$ denotes the layer normalization operater. Subsequently, $\boldsymbol{f}^{O}$ and $\hat{\boldsymbol{f}}^{I}$ are merged through the cross-attention mechanism, with the resulting fused features integrated back into the visual data stream through residual concatenation as
$\boldsymbol{f}^{I|O}=\boldsymbol{f}^{I}+\text{Cross}(\hat{\boldsymbol{f}}^{I},\boldsymbol{f}^{O}),\vspace{-1mm}$ (5)
@@
$\boldsymbol{f}^{V|O}=[\boldsymbol{f}^{I|O};\boldsymbol{f}^{H}],\vspace{-1mm}$ (6)
where [;][\ ;\ ][ ; ] represents the concatenation of two vectors. The feature $\boldsymbol{f}^{V|O}$ that enriches the visual information of the indicated object is fed to the subsequent module.
In this module, we counter-intuitively use the image as the query and the text as the key and value. This approach is driven by two considerations. First, it ensures that visual information is adequately preserved. Typically, instruction tokens are much shorter than image tokens, so a text-based query might not carry sufficient information. Second, the LDM (Rombach et al. 2022) conditions the image feature delivery on the text, also using the image as the query and the text as the key and value, indicating that such modal fusion is feasible. Our subsequent experiments confirm the effectiveness of this role assignment.
Object-focused Visual Semantic Learning
During the abstract concept learning phase, the original parameters of the image and text encoders are frozen, and the parameters of the AG-Adapter are trained. In each training batch, we sample $b$ image-text-object triples $\{\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{T}_{k},\boldsymbol{x}^{O}_{k}\}^{b}_{k=1}$ and encode them to obtain image-text feature pairs $\{\boldsymbol{y}^{I|O}_{k},\boldsymbol{y}^{T}_{k}\}^{b}_{k=1}$. Our goal is to train the AG-Adapter to extract the most relevant visual representations based on semantic instructions. As shown in Figure 3, to achieve this goal, we jointly optimize three training objectives that share model parameters.
@@
$\vspace{-1mm}s^{I|O}_{i,j}=s^{T}_{i,j}=(\boldsymbol{y}^{I|O}_{i})^{\top}\boldsymbol{y}^{T}_{j},\vspace{-1mm}$ (7)
where $s^{I|O}_{i,j}$ denotes the similarity of the $i$-th conditional image feature to the $j$-th text feature, and $s^{T}_{i,j}$ represents the similarity between the $i$-th text feature and the $j$-th text feature. For the conditional image feature $\boldsymbol{y}^{I|O}_{k}$ in feature pair $\{\boldsymbol{y}^{I|O}_{k},\boldsymbol{y}^{T}_{k}\}$, the corresponding text features $\boldsymbol{y}^{T}_{j|\boldsymbol{x}^{I}_{j}=\boldsymbol{x}^{I}_{k}\land\boldsymbol{x}^{O}_{j}=\boldsymbol{x}^{O}_{k}}$ of the same object within the same image are positive, while other text features within the batch are negative. The loss of a batch can be represented by
@@
$-\frac{1}{b}\log\frac{\sum_{k,j}\exp(s^{T}_{k,j}\mid\boldsymbol{x}^{I}_{j}=\boldsymbol{x}^{I}_{k}\land\boldsymbol{x}^{O}_{j}=\boldsymbol{x}^{O}_{k})}{\sum_{j}\exp(s^{T}_{k,j})},$ (9)
@@
Improv. - 41.2% - 39.8%
Table 1: Top-1 F1 and AUC (%) of zero-shot image classification on multi-object datasets. “Instruct.” signifies that classification is based on the instruction text rather than the image. “GroupViTy” and “GroupViTr”, i.e., GroupViT-gcc-yfcc and GroupViT-gcc-redcaps, denote two variants of GroupViT trained with different datasets, respectively. Since our evaluation metric accounts for and subtracts the influence of textual interference, it can result in negative values.
Model # TP Image $\rightarrow$ Text Text $\rightarrow$ Image
@@
Table 2: Fine-tuned image-text retrieval results on MOInst dataset. “# TP” denotes the number of trainable parameters.
Object-focused Image Discrimination (OID) is a binary classification task that requires the model to predict whether a given image and the indicated object match. For each batch of sample pairs $\{\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{O}_{k}\}^{b}_{k=1}$, we additionally construct $b$ negative pairs $\{\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{NO}_{k}\}^{b}_{k=1}$, where $\boldsymbol{x}^{NO}_{k}$ indicating object not appear in $\boldsymbol{x}^{I}_{k}$. These positive and negative samples $\{\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{O\cup NO}_{k}\}^{2b}_{k=1}$ are encoded to $\{\boldsymbol{y}^{I|O\cup NO}_{k}\}^{2b}_{k=1}$, which are then input into a binary linear classifier to obtain logits $\{\boldsymbol{z}_{k}\}^{2b}_{k=1}$. The loss function is formalized as
@@
$\mathcal{L}^{\text{OID}}$ $=-\frac{1}{2b}\sum_{i=1}^{2b}[t_{i}\log(p_{i})+(1-t_{i})\log(1-p_{i})],$ (14)
@@
Please refer to Appendix Figure 5 and the accompanying introduction for the inference pipeline.
Instruction of New Dataset
Although the Visual Genome (VG) dataset (Krishna et al. 2017) provides multiple objects and captions per image, it lacks object-caption correspondence, features noisy labels, and contains low-quality text. Therefore, it is not recommended to use the VG dataset directly. To address these issues, we construct the Multi-Object Instruction (MOInst) dataset. This dataset comprises 81,536 high-fidelity, complex images, accompanied by 244,378 textual captions. Each caption is associated with one of 264 distinct categories. Additional details are available in the Appendix.
Model # TP Inst. Img Enc. Type GiVE Post Proc. MOInst LVISf LVISa

### v2

@@ Abstract
Index Terms:
adapter, image encoder, instruction, multimodal learning, visual perception.
I Introduction
Multimodal Large Language Models (MLLMs) [1] have advanced general artificial intelligence with strong generation and inference capabilities in applications such as text-to-video generation [2], visual question answering [3], and embodied robotics [4]. A common architecture combines a visual encoder with a Large Language Model (LLM), embedding non-textual data into vectors interpretable by the LLM via a mapping mechanism. While research [5] highlights the effectiveness of this design, the quality of image embeddings from the visual encoder remains critical to MLLM performance.
An image encoder is a specialized visual encoder designed to map high-dimensional image data to a lower-dimensional feature space. These encoders can be broadly categorized based on their pre-training tasks into two main types: reconstruction-based and cross-modal contrastive learning-based encoders. Image encoding models trained with reconstruction tasks [6, 7] are proficient in capturing comprehensive image details. However, these models lack semantic alignment with text during training, which complicates the LLM’s ability to interpret image embeddings [8, 9]. Consequently, such encoders are infrequently utilized in MLLMs. Vision Transformer (ViT) models trained with image-text contrastive learning [10, 11] generally align effectively with LLMs but face an implicit “ignore” problem, limiting their expressive capability. This limitation arises because different modalities convey distinct types of information. For instance, an image may feature multiple objects with unique attributes, such as texture, color, spatial location, and potential interactions. In contrast, abstract text typically highlights only the most salient objects and provides limited descriptions of other visual elements. ViTs trained for image-text matching tend to focus on the salient regions of the image that correspond to the text, thereby overlooking secondary elements like the background. MLLMs using such visual encoders exhibit diminished response quality when users inquire about non-salient objects. In summary, for effective integration with LLMs, MLLMs require an image encoder that is both (1) semantically aligned with text during training and (2) capable of flexible attention to prevent the omission of relevant visual features.
@@ I Introduction
To address these challenges, we propose this Guiding Visual Encoder to perceive overlooked information (GiVE) approach, which aims to guide the visual encoder in adaptively adjusting its attention to well capture overlooked information. In this approach, we introduce a novel Attention-Guided Adapter (AG-Adapter) module that enhances the representation ability of the visual encoder by aligning the visual representations with abstract semantics. This module also functions as a plug-in for generalizing abstract semantics, enabling it to more effectively address user queries. To tackle the above limitations in detail, GiVE incorporates another innovative module: Object-focused Visual Semantic Learning. This module employs three distinct model loss terms: (i) an Object-focused Image-Text Contrast (OITC) loss, which encourages the model to generate distinct embeddings for varied instructions, ensuring consideration of both salient and non-salient objects; (ii) an Object-focused Image-Image Contrast (OIIC) loss, which improves the accuracy of object retrieval by enabling the model to learn common features among in-class objects, thus enhancing concept generalization ability; and (iii) an Object-focused Image Discrimination (OID) loss, which improves the comprehensiveness of object retrieval by facilitating the model in identifying specific objects and recognizing potential correlations among objects, thereby preventing the omission of objects. As illustrated in Fig. 1, GiVE enables visual encoders to generate image embeddings that are rich in targeted information.
@@
  • We propose an innovative GiVE approach to give attention to overlooked information, enabling dynamic adjustment of focus by modifying textual instructions. This capability allows for flexible and adaptive attention to various objects within the image.
@@
II GiVE
@@
II-A Model Architecture
@@
The overall architecture is depicted in Fig. 2. The proposed model includes a visual encoder $\Phi_{I}(\cdot,\cdot)$ and a text encoder $\Phi_{T}(\cdot)$ to respectively encode visual and textual content. It accepts image-text-object triplets $\{(\boldsymbol{x}^{I},\boldsymbol{x}^{T},\boldsymbol{x}^{O})\}$ as input, where $\boldsymbol{x}^{I}\in\mathcal{I}$ is an image, $\boldsymbol{x}^{T}\in\mathcal{T}$ is a text, and $\boldsymbol{x}^{O}\in\mathcal{O}$ is an indicative text denoting the target object, such as “person”. The model then extracts conditional image and text features $(\boldsymbol{y}^{I|O},\boldsymbol{y}^{T})$ using paired encoders. When extracting conditional image features, the instruction feature $\boldsymbol{y}^{O}$ is fused with the visual data stream within the AG-Adapter module. Formally,
$\boldsymbol{y}^{T}=\Phi_{T}(\boldsymbol{x}^{T}),\ \boldsymbol{y}^{O}=\Phi_{T}(\boldsymbol{x}^{O}),\ \boldsymbol{y}^{I|O}=\Phi_{I}(\boldsymbol{x}^{I},\boldsymbol{y}^{O}),\vspace{-1mm}$ (1)
@@
where $\boldsymbol{y}^{T}\in\mathbb{R}^{d}$ is text feature, $\boldsymbol{y}^{O}\in\mathbb{R}^{d}$ is instruction feature, and $\boldsymbol{y}^{I|O}$ is conditional image feature, i.e., the output of the visual encoder integrated with the AG-Adapter. The AG-Adapter is trained using our designed Object-focused Visual Semantic Learning component containing three loss terms: OITC, OIIC, and OID. Note that during the training phase, the loss is calculated based on the output embeddings of both encoders. During the inference phase, the text encoder is retained to serialize the textual instructions.
@@
Figure 3: Learning objectives illustration. The image and text encoders jointly compute three losses. (a) For Object-focused Image-Text Contrast, the paired text and image should not only correspond to each other but also correspond to the same semantic object. (b) Object-focused Image-Image Contrast requires the model to predict pairs of image features that contain the same semantic object. (c) Object-focused Image Discrimination determines whether a specific object exists in the image or not. The text instructions, such as “bike”, are pre-integrated into the prompt template designated as “a photo of {object}”.
II-B Attention-Guided Adapter
@@
The proposed AG-Adapter module, as highlighted in the green rectangle in Fig. 2, is a simple yet effective plug-in that interweaves semantic directives with visual cues, enabling the visual model to perceive queried objects. The AG-Adapter is inspired by the Latent Diffusion Model (LDM) [12], which enhances the alignment of visual representations with abstract semantics by conditioning on text representations. In particular, the AG-Adapter is incorporated into the pre-training feature extraction layers of the $\Phi_{I}(\cdot,\cdot)$, with only the inserted layer undergoing the training process.
@@
$\boldsymbol{f}^{V|O}=\varphi(\boldsymbol{y}^{O},\boldsymbol{f}^{V}),\vspace{-2mm}$ (2)
where $\boldsymbol{y}^{O}$ is instruction feature derived from user queries and $\boldsymbol{f}^{V}$ refers to the feature of the visual data flow. The $\boldsymbol{f}^{V}\in\mathbb{R}^{(M+K)\times d}$ is composed of two types of tokens: image tokens $\boldsymbol{f}^{I}\in\mathbb{R}^{M\times d}$ and other tokens $\boldsymbol{f}^{H}\in\mathbb{R}^{K\times d}$. The image tokens are derived directly from the input image, while the role of other tokens depends on the baseline model. In the case of GroupViT [13], other tokens are group tokens; in contrast, in CLIP [10], they are absent, and $K=0$. Within the AG-Adapter, the dual-layer MLP serves to bridge the gap between text and image representations, namely,
$\boldsymbol{f}^{O}=\text{MLP}(\boldsymbol{y}^{O}),\vspace{-3mm}$ (3)
@@
$\hat{\boldsymbol{f}}^{I}=\text{Norm}(\boldsymbol{f}^{I}),\vspace{-3mm}$ (4)
where $\text{Norm}(\cdot)$ denotes the layer normalization operator. Subsequently, $\boldsymbol{f}^{O}$ and $\hat{\boldsymbol{f}}^{I}$ are merged through the cross-attention mechanism, with the resulting fused features integrated back into the visual data stream through residual concatenation as
$\boldsymbol{f}^{I|O}=\boldsymbol{f}^{I}+\text{Cross}(\hat{\boldsymbol{f}}^{I},\boldsymbol{f}^{O}),\vspace{-3mm}$ (5)
@@
$\boldsymbol{f}^{V|O}=[\boldsymbol{f}^{I|O};\boldsymbol{f}^{H}],\vspace{-3mm}$ (6)
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] represents the concatenation of two vectors. The feature $\boldsymbol{f}^{V|O}$ that enriches the visual information of the indicated object is fed to the subsequent module.
In this module, we counter-intuitively use the image as the query and the text as the key and value. This approach is driven by two considerations. First, it ensures that visual information is adequately preserved. Typically, instruction tokens are much shorter than image tokens, so a text-based query might not carry sufficient information. Second, the LDM [12] conditions the image feature delivery on the text, also using the image as the query and the text as the key and value, indicating that such modal fusion is feasible. Our subsequent experiments confirm the effectiveness of this role assignment.
II-C Object-Focused Visual Semantic Learning
During the abstract concept learning phase, the original parameters of the image and text encoders are frozen, and the parameters of the AG-Adapter are trained. In each training batch, we sample $b$ image-text-object triples $\{(\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{T}_{k},\boldsymbol{x}^{O}_{k})\}^{b}_{k=1}$ and encode them to obtain image-text feature pairs $\{(\boldsymbol{y}^{I|O}_{k},\boldsymbol{y}^{T}_{k})\}^{b}_{k=1}$. Our goal is to train the AG-Adapter to extract the most relevant visual representations based on semantic instructions. As shown in Fig. 3, to achieve this goal, we jointly optimize three training objectives that share model parameters.
@@ II-C Object-Focused Visual Semantic Learning
$s^{I|O}_{i,j}=s^{T}_{i,j}=(\boldsymbol{y}^{I|O}_{i})^{\top}\boldsymbol{y}^{T}_{j},\vspace{-3mm}$ (7)
where $s^{I|O}_{i,j}$ denotes the similarity of the $i$-th conditional image feature to the $j$-th text feature, and $s^{T}_{i,j}$ represents the similarity between the $i$-th text feature and the $j$-th text feature. For the conditional image feature $\boldsymbol{y}^{I|O}_{k}$ in feature pair $\{(\boldsymbol{y}^{I|O}_{k},\boldsymbol{y}^{T}_{k})\}$, the corresponding text features $\boldsymbol{y}^{T}_{j|\boldsymbol{x}^{I}_{j}=\boldsymbol{x}^{I}_{k}\land\boldsymbol{x}^{O}_{j}=\boldsymbol{x}^{O}_{k}}$ of the same object within the same image are positive, while other text features within the batch are negative. The loss of a batch can be represented by
@@
$-\frac{1}{b}\log\frac{\sum_{k,j}\exp(s^{T}_{k,j}\mid\boldsymbol{x}^{I}_{j}=\boldsymbol{x}^{I}_{k}\land\boldsymbol{x}^{O}_{j}=\boldsymbol{x}^{O}_{k})}{\sum_{j}\exp(s^{T}_{k,j})},\vspace{-1mm}$ (9)
@@
TABLE I: Top-1 F1 and AUC (%) of zero-shot image classification on multi-object datasets
@@
Improv. 41.2% 39.8%
  • 1
    Classification is based on the instruction text rather than the image.
  • 2,3
    GroupViT-gcc-yfcc and GroupViT-gcc-redcaps, denote two variants of GroupViT trained with different datasets, respectively.
  • 4
    Since our evaluation metric accounts for and subtracts the influence of textual interference, it can result in negative values.
TABLE II: Fine-tuned image-text retrieval results on MOInst dataset
Model #Param.11footnotemark: 1 Image $\rightarrow$ Text Text $\rightarrow$ Image
@@
  • 1
    The number of trainable parameters.
Object-focused Image Discrimination (OID) is a binary classification task that requires the model to predict whether a given image and the indicated object match. For each batch of sample pairs $\{(\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{O}_{k})\}^{b}_{k=1}$, we additionally construct $b$ negative pairs $\{(\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{NO}_{k})\}^{b}_{k=1}$, where $\boldsymbol{x}^{NO}_{k}$ indicating object not appear in $\boldsymbol{x}^{I}_{k}$. These positive and negative samples $\{(\boldsymbol{x}^{I}_{k},\boldsymbol{x}^{O\cup NO}_{k})\}^{2b}_{k=1}$ are encoded to $\{\boldsymbol{y}^{I|O\cup NO}_{k}\}^{2b}_{k=1}$, which are then input into a binary linear classifier to obtain logits $\{\boldsymbol{z}_{k}\}^{2b}_{k=1}$. The loss function is formalized as
@@
$\mathcal{L}^{\text{OID}}$ $=-\frac{1}{2b}\sum_{i=1}^{2b}[t_{i}\log(p_{i})+(1-t_{i})\log(1-p_{i})],\vspace{-1mm}$ (14)
@@
II-D Instruction of New Dataset
Although the Visual Genome (VG) dataset [14] provides multiple objects and captions per image, it lacks object-caption correspondence, features noisy labels, and contains low-quality text. Therefore, it is not recommended to use the VG dataset directly. To address these issues, we construct the Multi-Object Instruction (MOInst) dataset. This dataset comprises 81,536 high-fidelity, complex images, accompanied by 244,378 textual captions. Each caption is associated with one of 264 distinct categories.
TABLE III: The effect of semantic instructions on secondary coding methods. AG-CLIP is CLIP-ViT with GiVE for attention guidance
Model #Param. Inst. Img Enc. Type GiVE Post Proc. MOInst LVISf LVISa